# Strand, London WC2R 2LS, UK

### v1

@@ A.2 Variation of the Gauss-Bonnet Action
Suppose now that we add in (93) a Gauss-Bonnet term:
$\mathcal{I}_{\text{GB}}=-\frac{\alpha}{(d-2)(d-3)16\pi G_{N}}\int_{\mathcal{M}}d^{d+1}x\sqrt{g}\left(R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}-4R^{\mu\nu}R_{\mu\nu}+R^{2}\right)\ .$ (115)
Recall that the coupling $\alpha$ is dimensionful and appropriately normalized by a factor of $1/32\pi^{2}$ so that $\mathcal{I}_{\text{GB}}$ reduces to the Euler character in $4$-dimensions. The variation of the Gauss-Bonnet term turns out to be Bunch1981 ; Myers:1987yn ; Davis:2002gn ; Deruelle:2017xel : [^29] We again ignore corner terms, for a treatment see ref. Chakraborty:2017zep .
$\frac{(d-2)(d-3)16\pi G_{N}}{\alpha}\delta\mathcal{I}_{\text{GB}}=-\int_{\mathcal{M}}d^{d+1}x\sqrt{g}\ \mathcal{E}_{\mu\nu}\delta g^{\mu\nu}$ $-\epsilon\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\ \mathcal{B}_{\mu\nu}\delta h^{\mu\nu}$ (116)
$+\epsilon\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\ \mathcal{Q}\right)\ ,$
@@
We can again impose Dirichlet boundary conditions $\delta h_{\mu\nu}|_{\partial\mathcal{M}}=0$, which kill the second term in eq. (116) and therefore the action has to be supplemented by a Bunch-Myers (BM) term Bunch1981 ; Myers:1987yn :
$\mathcal{I}_{\text{BM}}=-\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\ .$ (119)
As in GR, the coefficient of this boundary term is the same irrespective of dimension.
@@
Merging eqs. (94) and (116) one observes that fixing the combination:
@@
As in the instance of pure GR, one could try to fix the corresponding canonically conjugate momentum, which in the GB case reads:
$\not{\mathcal{B}}^{\mu\nu}:=\frac{\delta}{\delta h_{\mu\nu}}\left(\mathcal{I}_{\text{GB}}+\mathcal{I}_{\text{BM}}\right)=-\frac{\epsilon\alpha\sqrt{h}}{(d-2)(d-3)16\pi G_{N}}\mathcal{B}^{\mu\nu}\ .$ (121)
Therefore, eq. (116) becomes ($\mathcal{B}:=h^{\mu\nu}\mathcal{B}_{\mu\nu}$):
$\delta\mathcal{I}_{\text{GB}}-\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)$ $=\epsilon\int_{\partial\mathcal{M}}d^{d}y\not{\mathcal{B}}^{\mu\nu}\delta h_{\mu\nu}$ (122)
$=\epsilon\delta\int_{\partial\mathcal{M}}d^{d}y\not{\mathcal{B}}-\epsilon\int_{\partial\mathcal{M}}d^{d}yh_{\mu\nu}\delta\not{\mathcal{B}}^{\mu\nu}\ ,$
@@
Consequently, the trace of the GB momentum is:
$\not{\mathcal{B}}=\frac{\epsilon\alpha\sqrt{h}}{(d-2)32\pi G_{N}}\mathcal{Q}\ ,$ (125)
and thus the variation (122) becomes on-shell:
$\delta\mathcal{I}_{\text{GB}}-\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\hat{=}$ $\frac{\epsilon\alpha}{(2-d)32\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)$ (126)
$+\epsilon\int_{\partial\mathcal{M}}d^{d}yh_{\mu\nu}\delta\not{\mathcal{B}}^{\mu\nu}\ ,$
@@
$\delta\mathcal{I}_{\text{GB}}\hat{=}\frac{\epsilon(5-d)\alpha}{(d-2)(d-3)32\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\ ,$ (127)
@@
$\mathcal{I}_{\text{NBCs}}^{(\text{GB})}=\frac{\epsilon(d-5)\alpha}{(d-2)(d-3)32\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\ .$ (128)
The boundary term does not appear in $6$-dimensions, as did the York term for NBCs in GR (100) in $4$-dimensions. This is a general pattern for Lovelock gravity theories, namely the coefficient of the boundary term for the generalized Neumann boundary conditions vanishes in $(2n+2)$-dimensions, when the Lovelock action contains the Euler density of $2n$-dimensions.
@@
Focusing on the conformal case, the second term of the variation (116) becomes (using eq. (102)):
$\int_{\partial\mathcal{M}}d^{d}y\ \not{\mathcal{B}}^{\mu\nu}\delta h_{\mu\nu}$ $=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\not{\mathcal{B}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\int_{\partial\mathcal{M}}d^{d}y\frac{2}{d\sqrt{h}}\not{\mathcal{B}}\delta\sqrt{h}$ (129)
$=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\not{\mathcal{B}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{\alpha}{d(d-2)16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\ \mathcal{Q}\delta\sqrt{h}$
$=\int_{\partial\mathcal{M}}d^{d}y$ $\ h^{\frac{1}{d}}\not{\mathcal{B}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{\alpha}{d(d-2)16\pi G_{N}}\left[\delta\left(\int_{\partial\mathcal{M}}d^{d}y\ \mathcal{Q}\sqrt{h}\right)-\int_{\partial\mathcal{M}}d^{d}y\ \sqrt{h}\delta\mathcal{Q}\right]\ .$
Consequently, from eq. (122) we have (omitting for a moment the terms fixed by the boundary condition $\delta\tilde{h}_{\mu\nu}|_{\partial\mathcal{M}}=0$):
$\delta\mathcal{I}_{\text{EGB}}-\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\hat{=}-\frac{\epsilon\alpha}{d(d-2)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\ \mathcal{Q}\sqrt{h}\right)\ ,$ (130)
@@
$\delta\mathcal{I}_{\text{GB}}\hat{=}\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\ .$ (131)
Re-instating all the terms:
$\delta\mathcal{I}_{\text{EGB}}\hat{=}-\int_{\partial\mathcal{M}}d^{d}y\tilde{\not{\mathcal{B}}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}-\int_{\partial\mathcal{M}}d^{d}y\ \pi_{\mathcal{Q}}\delta\mathcal{Q}+\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\ ,$ (132)
where we defined:
$\begin{cases}\tilde{\not{\mathcal{B}}}^{\mu\nu}:=h^{\frac{1}{d}}\left(\not{\mathcal{B}}^{\mu\nu}-\dfrac{1}{d}h^{\mu\nu}\not{\mathcal{B}}\right)\ ,\\ \pi_{\mathcal{Q}}:=-\dfrac{\epsilon\alpha\sqrt{h}}{d(d-2)16\pi G_{N}}\ .\end{cases}$ (133)
Therefore, the overall on-shell variation of the EH and GB actions reads:
$\delta\mathcal{I}_{\text{EGB}}$ $\hat{=}-\int_{\partial\mathcal{M}}d^{d}y\left(\tilde{\pi}^{\mu\nu}+\tilde{\not{\mathcal{B}}}^{\mu\nu}\right)\delta\tilde{h}_{\mu\nu}-\int_{\partial\mathcal{M}}d^{d}y\ \left(\pi_{\mathcal{K}}\delta\mathcal{K}+\pi_{\mathcal{Q}}\delta\mathcal{Q}\right)$ (134)
$+\frac{\epsilon}{d8\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)+\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\ ,$
and using eqs. (105), (133):
$\delta\mathcal{I}_{\text{EGB}}$ $\hat{=}-\int_{\partial\mathcal{M}}d^{d}y\left(\tilde{\pi}^{\mu\nu}+\tilde{\not{\mathcal{B}}}^{\mu\nu}\right)\delta\tilde{h}_{\mu\nu}+\frac{\epsilon}{d16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\ \left(2(d-1)\delta\mathcal{K}+\frac{\alpha}{d-2}\delta\mathcal{Q}\right)$ (135)
$+\frac{\epsilon}{d8\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{K}\right)+\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\right)\ .$
This last relation implies that we have to modify the GR conformal boundary conditions (3), where we again fix the conformal class of the induced metric along the boundary but now we fix the following scalar quantity:
@@
leading to the ECBC (12). Then, apart from the York term of CBC, one has to further add the following boundary term
$\mathcal{I}_{\text{ECBC}}=-\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\ ,$ (138)
@@
There exists a GB generalization of the one parameter family of GBCs of ref. Liu:2024ymn . In particular, plugging an arbitrary coefficient $\Phi$ in front of the BM term
$\mathcal{I}_{\text{GCBC}}^{(\text{GB})}=-\frac{\Phi}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\mathcal{Q}\ ,$ (139)
@@
$\delta\mathcal{I}_{\text{EGB}}\hat{=}\frac{\epsilon}{16\pi G_{N}}$ $\int_{\partial\mathcal{M}}d^{d}y\Bigg[\tilde{\Pi}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+2\left(1-\Theta-\frac{d-1}{d}\right)\mathcal{K}\delta\sqrt{h}+2(1-\Theta)\sqrt{h}\delta\mathcal{K}$ (140)
$+\frac{\alpha}{(d-2)(d-3)}\left(\left(1-\Phi-\frac{d-3}{d}\right)\mathcal{Q}\delta\sqrt{h}+(1-\Phi)\sqrt{h}\delta\mathcal{Q}\right)\Bigg]\ ,$
@@
$\delta\mathcal{I}_{\text{EGB}}\hat{=}\frac{1-\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}h^{-p}\delta(h^{p}\mathcal{P})\ ,$ (142)
again with $p=\frac{1}{2(1-\Theta)}(1-\Theta-\frac{d-1}{d})$. Therefore, the extended generalized boundary conditions (EGBCs) read
@@
with the total boundary term supplementing the EGB bulk action being
$\mathcal{I}_{\text{EGBCs}}=-\frac{1}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{h}\left(2\Theta\mathcal{K}+\frac{\alpha}{(d-2)(d-3)}\Phi\mathcal{Q}\right)\ ,$ (145)
@@
For $p=0$ one recovers the ECBC (12), for $p\rightarrow\infty$ the Dirichlet York (96) and BM (119) terms, while in the limit $\alpha\rightarrow 0$ one goes back to the boundary conditions (113) and (114).

### v2

@@ A.2 Variation of the Gauss-Bonnet Action
Suppose now that we add in (93) a Gauss-Bonnet term,
$\mathcal{I}_{\text{GB}}=\frac{\alpha}{(d-2)(d-3)16\pi G_{N}}\int_{\mathcal{M}}d^{d+1}x\sqrt{-g}\left(R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}-4R^{\mu\nu}R_{\mu\nu}+R^{2}\right)\ .$ (114)
Recall that the coupling $\alpha$ is dimensionful and appropriately normalized by a factor of $1/32\pi^{2}$ so that $\mathcal{I}_{\text{GB}}$ reduces to the Euler character in $4$-dimensions. The variation of the Gauss-Bonnet term turns out to be Bunch (1981); Myers (1987a); Davis (2003); Deruelle et al. (2018), [^30] We again ignore corner terms, for a treatment see ref. Chakraborty et al. (2017).
$\frac{(d-2)(d-3)16\pi G_{N}}{\alpha}\delta\mathcal{I}_{\text{GB}}=\int_{\mathcal{M}}d^{d+1}x\sqrt{-g}\ \mathcal{E}_{\mu\nu}\delta g^{\mu\nu}$ $+\epsilon\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\ \mathcal{B}_{\mu\nu}\delta h^{\mu\nu}$ (115)
$-\epsilon\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\ \mathcal{Q}\right)\ ,$
@@
We can again impose Dirichlet boundary conditions $\delta h_{\mu\nu}|_{\partial\mathcal{M}}=0$, such that the action has to be supplemented by the Bunch-Myers (BM) term Bunch (1981); Myers (1987a),
$\mathcal{I}_{\text{BM}}=\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\ .$ (118)
As in GR, the coefficient of this boundary term is the same irrespective of dimension, apart from the particular dimension-dependent normalization of the GB coupling we have chosen.
@@
Merging eqs. (94) and (115) one observes that fixing the combination
@@
As in the instance of pure GR, one could try to fix the corresponding canonically conjugate momentum, which in the GB case reads
$\not{\mathcal{B}}^{\mu\nu}:=\frac{\delta}{\delta h_{\mu\nu}}\left(\mathcal{I}_{\text{GB}}+\mathcal{I}_{\text{BM}}\right)=-\frac{\epsilon\alpha\sqrt{|h|}}{(d-2)(d-3)16\pi G_{N}}\mathcal{B}^{\mu\nu}\ .$ (120)
Therefore, eq. (115) becomes ($\mathcal{B}:=h^{\mu\nu}\mathcal{B}_{\mu\nu}$),
$\delta\mathcal{I}_{\text{GB}}+\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)$ $=-\epsilon\int_{\partial\mathcal{M}}d^{d}y\not{\mathcal{B}}^{\mu\nu}\delta h_{\mu\nu}$ (121)
$=-\epsilon\delta\int_{\partial\mathcal{M}}d^{d}y\not{\mathcal{B}}+\epsilon\int_{\partial\mathcal{M}}d^{d}yh_{\mu\nu}\delta\not{\mathcal{B}}^{\mu\nu}\ ,$
@@
Consequently, the trace of the GB momentum is,
$\not{\mathcal{B}}=\frac{\epsilon\alpha\sqrt{|h|}}{(d-2)32\pi G_{N}}\mathcal{Q}\ ,$ (124)
and thus the on-shell variation (121) becomes,
$\delta\mathcal{I}_{\text{GB}}+\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\hat{=}$ $-\frac{\epsilon\alpha}{(2-d)32\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)$ (125)
$-\epsilon\int_{\partial\mathcal{M}}d^{d}yh_{\mu\nu}\delta\not{\mathcal{B}}^{\mu\nu}\ .$
@@
$\delta\mathcal{I}_{\text{GB}}\hat{=}-\frac{\epsilon(5-d)\alpha}{(d-2)(d-3)32\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\ ,$ (126)
@@
$\mathcal{I}_{\text{NBCs}}^{(\text{GB})}=-\frac{\epsilon(d-5)\alpha}{(d-2)(d-3)32\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\ .$ (127)
The boundary term does not appear in $d=5$, as did the York term for NBCs in GR (100) in $d=3$. This is a general pattern for Lovelock gravity theories, namely the coefficient of the boundary term for the generalized Neumann boundary conditions vanishes in $d=2n+1,\,n\in\mathbb{N}^{\star}$, when the Lovelock action contains the Euler density of d=2n1d=2n-1.
@@
Focusing on the conformal case, the second term of the variation (115) becomes (using eq. (102)),
$\int_{\partial\mathcal{M}}d^{d}y\ \not{\mathcal{B}}^{\mu\nu}\delta h_{\mu\nu}$ $=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\not{\mathcal{B}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\int_{\partial\mathcal{M}}d^{d}y\frac{2}{d\sqrt{|h|}}\not{\mathcal{B}}\delta\sqrt{|h|}$ (128)
$=\int_{\partial\mathcal{M}}d^{d}y\ h^{\frac{1}{d}}\not{\mathcal{B}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{\alpha}{d(d-2)16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\ \mathcal{Q}\delta\sqrt{|h|}$
$=\int_{\partial\mathcal{M}}d^{d}y$ $\ h^{\frac{1}{d}}\not{\mathcal{B}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\frac{\alpha}{d(d-2)16\pi G_{N}}\left[\delta\left(\int_{\partial\mathcal{M}}d^{d}y\ \mathcal{Q}\sqrt{|h|}\right)-\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\delta\mathcal{Q}\right]\ .$
Consequently, from eq. (121) we have (omitting for a moment the terms fixed by the boundary condition $\delta\tilde{h}_{\mu\nu}|_{\partial\mathcal{M}}=0$),
$\delta\mathcal{I}_{\text{EGB}}+\frac{\epsilon\alpha}{(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\hat{=}\frac{\epsilon\alpha}{d(d-2)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\ \mathcal{Q}\sqrt{|h|}\right)\ ,$ (129)
@@
$\delta\mathcal{I}_{\text{GB}}\hat{=}-\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\ .$ (130)
Re-instating all the terms
$\delta\mathcal{I}_{\text{EGB}}\hat{=}\int_{\partial\mathcal{M}}d^{d}y\tilde{\not{\mathcal{B}}}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+\int_{\partial\mathcal{M}}d^{d}y\ \pi_{\mathcal{Q}}\delta\mathcal{Q}-\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\ ,$ (131)
where we defined
$\begin{cases}\tilde{\not{\mathcal{B}}}^{\mu\nu}:=h^{\frac{1}{d}}\left(\not{\mathcal{B}}^{\mu\nu}-\dfrac{1}{d}h^{\mu\nu}\not{\mathcal{B}}\right)\ ,\\ \pi_{\mathcal{Q}}:=-\dfrac{\epsilon\alpha\sqrt{|h|}}{d(d-2)16\pi G_{N}}\ .\end{cases}$ (132)
Therefore, the overall on-shell variation of the EH and GB actions reads,
$\delta\mathcal{I}_{\text{EGB}}$ $\hat{=}\int_{\partial\mathcal{M}}d^{d}y\left(\tilde{\pi}^{\mu\nu}+\tilde{\not{\mathcal{B}}}^{\mu\nu}\right)\delta\tilde{h}_{\mu\nu}+\int_{\partial\mathcal{M}}d^{d}y\ \left(\pi_{\mathcal{K}}\delta\mathcal{K}+\pi_{\mathcal{Q}}\delta\mathcal{Q}\right)$ (133)
$-\frac{\epsilon}{d8\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)-\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\ ,$
and using eqs. (105), (132),
$\delta\mathcal{I}_{\text{EGB}}\hat{=}$ $\int_{\partial\mathcal{M}}d^{d}y\left(\tilde{\pi}^{\mu\nu}-\tilde{\not{\mathcal{B}}}^{\mu\nu}\right)\delta\tilde{h}_{\mu\nu}-\frac{\epsilon}{d16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\ \left(2(d-1)\delta\mathcal{K}+\frac{\alpha}{d-2}\delta\mathcal{Q}\right)$ (134)
$-\frac{\epsilon}{d8\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{K}\right)-\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\delta\left(\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\right)\ .$
This last relation implies that we have to modify the GR conformal boundary conditions (3), where we again fix the conformal class of the induced metric along the boundary but now we fix the following scalar quantity,
@@
leading to the ECBC (12). Then, apart from the boundary term for CBC, one has to further add the following boundary term
$\mathcal{I}_{\text{ECBC}}=\frac{3\epsilon\alpha}{d(d-2)(d-3)16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\ ,$ (137)
@@
There exists a GB generalization of the one parameter family of GBCs of ref. Liu et al. (2024). In particular, plugging an arbitrary coefficient $\Phi$ in front of the BM term
$\mathcal{I}_{\text{GCBC}}^{(\text{GB})}=\frac{\epsilon\Phi}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\mathcal{Q}\ ,$ (138)
@@
$\delta\mathcal{I}_{\text{EGB}}\hat{=}-\frac{\epsilon}{16\pi G_{N}}$ $\int_{\partial\mathcal{M}}d^{d}y\Bigg[\tilde{\Pi}^{\mu\nu}\delta\tilde{h}_{\mu\nu}+2\left(1-\Theta-\frac{d-1}{d}\right)\mathcal{K}\delta\sqrt{|h|}+2(1-\Theta)\sqrt{|h|}\delta\mathcal{K}$ (139)
$+\frac{\alpha}{(d-2)(d-3)}\left(\left(1-\Phi-\frac{d-3}{d}\right)\mathcal{Q}\delta\sqrt{|h|}+(1-\Phi)\sqrt{|h|}\delta\mathcal{Q}\right)\Bigg]\ ,$
@@
$\delta\mathcal{I}_{\text{EGB}}\hat{=}-\epsilon\frac{1-\Theta}{8\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}h^{-p}\delta(h^{p}\mathcal{P})\ ,$ (141)
again with $p=\frac{1}{2(1-\Theta)}(1-\Theta-\frac{d-1}{d})$. Therefore, the extended generalized boundary conditions (EGBCs) have the following quantities fixed,
@@
when the total boundary term supplementing the EGB bulk action is given by
$\mathcal{I}_{\text{EGBCs}}=\frac{\epsilon}{16\pi G_{N}}\int_{\partial\mathcal{M}}d^{d}y\sqrt{|h|}\left(2\Theta\mathcal{K}+\frac{\alpha}{(d-2)(d-3)}\Phi\mathcal{Q}\right)\ ,$ (144)
@@
For $p=0$ one recovers the ECBC (12); for $p\rightarrow\infty$, the Dirichlet  (96) and (118) terms. In the limit $\alpha\rightarrow 0$ one goes back to the boundary conditions (112) and (113).